# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v1

@@ B.1 Simplicial Complexes
A $d$-dimensional simplex is formed by a set of $(d+1)$ interacting nodes and includes all the subsets of $\delta+1$ nodes (with $\delta<d$), which are called the $\delta$-dimensional faces of the simplex. A simplicial complex of dimension $d$ is formed by simplices of dimension at most equal to $d$ glued along their faces.
Definition B.1 (Simplicial complexes).
@@
B.2 Cell Complexes
Cell complexes generalize simplicial complexes by incorporating generalized building blocks called cells instead of relying solely on simplices (Hatcher, 2001). This broader approach allows for the representation of many-body interactions that do not adhere to the strict requirements of simplicial complexes. For example, a square can be interpreted as a cell of four-body interactions whose faces are just four links. This flexibility is advantageous in scenarios such as social networks, where, for instance, a discussion group might not involve all-to-all pairwise interactions, or in protein interaction networks, where proteins in a complex may not bind pairwise.
@@ B.2 Cell Complexes
As shown in Fig. 5 a and b, triangles and tetrahedrons are two particular types of cell complexes called simplicial complexes (SCs). The only 2-cells they allow are triangle-shaped. The sphere shown in Fig. 5 c is a 2-dimensional cell complex. It is constructed using two 0-cells (i.e., nodes), connected by two 1-cells (i.e., the edges forming the equator). The equator serves as the boundary for two 2-dimensional disks (the hemispheres), which are glued together along the equator to form the sphere. The torus in Fig. 5 d is a 2-dimensional cell complex formed by attaching a single 1-cell to itself in two directions to form the loops of the torus. The resulting structure is then completed by attaching a 2-dimensional disk, forming the surface of the torus. Note that this is just one way to represent the torus as a cell complex, and other decompositions might lead to different numbers of cells and faces.
Appendix C Additional Explanation on Related Works
C.1 Graph Generative Models
@@
Graph generation has been extensively studied, which dates back to the early works of the random network models, such as the Erdős–Rényi (ER) model (Erdős et al., 1960) and the Barabási-Albert (BA) model (Barabási & Albert, 1999). Recent graph generative models make great progress in graph distribution learning by exploiting the capacity of deep neural networks. GraphRNN (You et al., 2018) and GraphVAE (Simonovsky & Komodakis, 2018) adopt sequential strategies to generate nodes and edges. MolGAN (De Cao & Kipf, 2018) integrates generative adversarial networks (GANs) with reinforcement learning objectives to synthesize molecules with desired chemical properties. Shi et al. (2020) generates molecular graphs using a flow-based approach, while GraphDF (Luo et al., 2021) adopts an autoregressive flow-based model with discrete latent variables. Additionally, GraphEBM (Liu et al., 2021) employs an energy-based model for molecular graph generation. However, the end-to-end structure of these methods often makes them more challenging to train compared to diffusion-based generative models.
C.2 Diffusion-based Generative Models
@@
A leap in graph generative models has been marked by the recent progress in diffusion-based generative models (Song et al., 2021). EDP-GNN (Niu et al., 2020) generates the adjacency matrix by learning the score function of the denoising diffusion process, while GDSS (Jo et al., 2022) extends this framework by simultaneously generating node features and an adjacency matrix with a joint score function capturing the node-edge dependency. DiGress (Vignac et al., 2023) addresses the discretization challenge due to Gaussian noise, while CDGS (Huang et al., 2023) designs a conditional diffusion model based on discrete graph structures. GSDM (Luo et al., 2023a) introduces an efficient graph diffusion model driven by low-rank diffusion SDEs on the spectrum of adjacency matrices. GPrinFlowNet (Mo et al., 2024) proposes a semantic-preserving framework based on a low-to-high frequency generation curriculum, where the $k$-th intermediate generation state corresponds to the $k$ smallest principal components of the adjacency matrices. Despite these advancements, current methods are ineffective at modelling the topological properties of higher-order systems since learning to denoise the noisy samples does not explicitly lead to preserving the intricate structural dependencies required for generating realistic graphs.
C.3 Diffusion Bridge
@@
Several recent works have improved the generative framework of diffusion models by leveraging the diffusion bridge processes, i.e., processes conditioned to the endpoints. Wu et al. (2022) inject physical information into the process by incorporating informative prior to the drift. GLAD (Boget et al., 2024) employs the Brownian bridge on a discrete latent space with endpoints conditioned on data samples. GruM (Jo et al., 2024) utilizes the OU bridge to condition the diffusion endpoint as the weighted mean of all possible final graphs. However, existing methods often overlook or inadvertently disrupt the higher-order topological structures in the graph generation process.
Appendix D Details for Higher-order Guided Generation
D.1 Denoising Network Parametrization
The denoising network in HOG-Diff is a critical component responsible for estimating the score functions required to reverse the diffusion process effectively. The architecture of the proposed denoising network is depicted in Fig. 6. The input $\bm{A}_{t}$ is computed from $\bm{U}_{0}$ and $\bm{\Lambda}_{t}^{(k)}$ using the relation $\bm{A}_{t}=\bm{D}_{t}^{(k)}-\bm{L}^{(k)}_{t}$, where the Laplacian matrix is given by $\bm{L}^{(k)}_{t}=\bm{U}_{0}\bm{\Lambda}_{t}^{(k)}\bm{U}_{0}^{\top}$ and the diagonal degree matrix is given by $\bm{D}_{t}^{(k)}=\operatorname{diag}\left(\bm{L}_{t}^{(k)}\right)$. To enhance the input to the Attention module, we derive enriched node and edge features using the $l$-step random walk matrix obtained from the binarized $\bm{A}_{t}$. Specifically, the arrival probability vector is incorporated as additional node features, while the truncated shortest path distance derived from the same matrix is employed as edge features. Temporal information is integrated into the outputs of the Attention and GCN modules using Feature-wise Linear Modulation (FiLM) (Perez et al., 2018) layers, following sinusoidal position embeddings (Waswani et al., 2017).
Figure 6: Denoising Network Architecture of HOG-Diff. The denoising network integrates GCN and Attention blocks to capture both local and global features, and further incorporates time information through FiLM layers. These enriched outputs are subsequently concatenated and processed by separate feed-forward networks to produce predictions for $\nabla_{\bm{X}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}_{t}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, respectively.
A graph processing module is considered permutation invariant if its output remains unchanged under any permutation of its input, formally expressed as $f(\bm{G})=x\iff f(\pi(\bm{G}))=x$, where $\pi(\bm{G})$ represents a permutation of the input graph $\bm{G}$. It is permutation equivariant when the output undergoes the same permutation as the input, formally defined as $f(\pi(\bm{G}))=\pi(f(\bm{G}))$. It is worth noting that our denoising network model is permutation equivalent as each model component avoids any node ordering-dependent operations.
D.2 Training and Sampling Proceudre
@@
As shown in Fig. 7, HOG-Diff implements a coarse-to-fine generation curriculum, with the forward diffusion and reverse denoising processes divided into $K$ easy-to-learn subprocesses. Each subprocess is realized using the generalized OU bridge process.
@@
We provide the pseudo-code of the training process in Alg. 1. In our experiments, we adopt a two-step generation process, i.e., $K=2$. We initialize $\mathcal{S}^{(0)}=\bm{G}$; under this specific condition, the cell complex filtering operation returns the input unchanged. The set $\mathcal{S}^{(1)}$ corresponds to the 2-cell complex for molecule generation tasks or the 3-simplicial complex for generic graph generation tasks. We set $\mathcal{S}^{(2)}=\varnothing$, and for this particular case, we define the cell complex filtering function as $\operatorname{CCF}(\bm{G},\varnothing)=\mathcal{N}(\bm{0},\bm{I})$.

### v2

@@ B.1 Higher-order Networks
A $d$-dimensional simplex is formed by a set of $(d+1)$ interacting nodes and includes all the subsets of $\delta+1$ nodes (with $\delta<d$), which are called the $\delta$-dimensional faces of the simplex. A simplicial complex of dimension $d$ is formed by simplices of dimension at most $d$ glued along their faces.
Definition 6 (Simplicial complexes).
@@
Cell Complexes
Cell complexes (CCs) generalize simplicial complexes by incorporating generalized building blocks called cells instead of relying solely on simplices (Hatcher, 2001). This broader approach allows for the representation of many-body interactions that do not adhere to the strict requirements of simplicial complexes. For example, a square can be interpreted as a cell of four-body interactions whose faces are just four links. This flexibility is advantageous in scenarios such as social networks, where, for instance, a discussion group might not involve all-to-all pairwise interactions, or in protein interaction networks, where proteins in a complex may not bind pairwise.
@@
As shown in Fig. 5 (a) and (b), triangles and tetrahedra are two particular types of cell complexes called simplicial complexes (SCs). The only 2-cells they allow are triangle-shaped. The sphere shown in Fig. 5 (c) is a 2-dimensional cell complex. It is constructed using two 0-cells (i.e., nodes), connected by two 1-cells (i.e., the edges forming the equator). The equator serves as the boundary for two 2-dimensional disks (the hemispheres), which are glued together along the equator to form the sphere. The torus in Fig. 5 (d) is a 2-dimensional cell complex formed by attaching a single 1-cell to itself in two directions to form the loops of the torus. The resulting structure is then completed by attaching a 2-dimensional disk, forming the surface of the torus. Note that this is just one way to represent the torus as a cell complex, and other decompositions might lead to different numbers of cells and faces.
These topological frameworks provide the mathematical foundation for capturing multi-way interactions beyond pairwise graphs. Building upon this background, we next review advances in graph generative models, highlighting how existing approaches attempt to learn graph distributions and where they fall short in preserving such higher-order structures.
B.2 Graph Generative Models
The study of graph generation seeks to synthesize graphs that align with the observed distribution. Graph generation has been extensively studied, which dates back to the early works of the random network models, such as the Erdős–Rényi (ER) model (Erdős et al., 1960) and the Barabási-Albert (BA) model (Barabási & Albert, 1999). While these models offer foundational insights, they are often too simplistic to capture the complexity of graph distributions we encounter in practice.
@@ B.2 Graph Generative Models
Recent graph generative models have made great progress in graph distribution learning by exploiting the capacity of deep neural networks. GraphRNN (You et al., 2018) and GraphVAE (Simonovsky & Komodakis, 2018) adopt sequential strategies to generate nodes and edges. MolGAN (De Cao & Kipf, 2018) integrates generative adversarial networks (GANs) with reinforcement learning objectives to synthesize molecules with desired chemical properties. Shi et al. (2020) generates molecular graphs using a flow-based approach, while GraphDF (Luo et al., 2021) adopts an autoregressive flow-based model with discrete latent variables. Additionally, GraphEBM (Liu et al., 2021) employs an energy-based model for molecular graph generation. However, the end-to-end structure of these methods often makes them more challenging to train compared to diffusion-based generative models.
Diffusion-based Generative Models
@@
A leap in graph generative models has been marked by the recent progress in diffusion-based generative models (Song et al., 2021). EDP-GNN (Niu et al., 2020) generates the adjacency matrix by learning the score function of the denoising diffusion process, while GDSS (Jo et al., 2022) extends this framework by simultaneously generating node features and an adjacency matrix with a joint score function capturing the node-edge dependency. DiGress (Vignac et al., 2023) addresses the discretization challenge due to Gaussian noise, while CDGS (Huang et al., 2023) designs a conditional diffusion model based on discrete graph structures. GSDM (Luo et al., 2023a) introduces an efficient graph diffusion model driven by low-rank diffusion SDEs on the spectrum of adjacency matrices. HypDiff (Fu et al., 2024) introduces a geometrically latent diffusion on hyperbolic space to preserve the anisotropy of the graph. Despite these advancements, current methods are ineffective at modeling the topological properties of higher-order systems since learning to denoise the noisy samples does not explicitly lead to preserving the intricate structural dependencies required for generating realistic graphs.
Diffusion Bridge
@@
Diffusion bridge processes, i.e., processes conditioned to the endpoints, have been widely adopted in image-related domains, including image generation (De Bortoli et al., 2021), image translation (Zhou et al., 2024), and image restoration (Luo et al., 2023b; Yue et al., 2024). Recently, several studies have improved the graph generative framework of diffusion models by leveraging the diffusion bridge processes. Wu et al. (2022) inject physical information into the process by incorporating informative prior to the drift. GLAD (Boget et al., 2024) employs the Brownian bridge on a discrete latent space with endpoints conditioned on data samples. GruM (Jo et al., 2024) utilizes the OU bridge to condition the diffusion endpoint as the weighted mean of all possible final graphs. However, existing methods often overlook or inadvertently disrupt the higher-order topological structures in the graph generation process.
Hierarchical and Fragment-based Generation
Several recent studies have also explored hierarchical and fragment-based generative frameworks. HiGen (Karami, 2024) decomposes graph generation into multiple layers of abstraction, using separate neural networks to model intra-community structures and inter-community connections at each level. GPrinFlowNet (Mo et al., 2024) proposes a semantic-preserving framework based on a low-to-high frequency generation curriculum, where the $k$-th intermediate generation state corresponds to the $k$ smallest principal components of the adjacency matrices. Dymond (Zeno et al., 2021) focuses on temporal motifs in dynamic graph generation. HierDiff (Qiang et al., 2023) progressively generates fragment-level 3D geometries, refines them into fine-grained fragments, and then assembles these fragments into complete molecules. MiCaM (Geng et al., 2023) synthesizes molecules by iteratively merging motifs.
Higher-order Generative Models
Since higher-order structures are intrinsic to many real-world systems, incorporating them into generative models could yield more faithful representations of complex phenomena. Existing efforts have primarily explored higher-order information through hypergraphs. HypeBoy (Kim et al., 2024) introduces a self-supervised hypergraph representation framework based on a hyperedge filling task, which enhances embeddings rather than performing direct generation. Hygene (Gailhard et al., 2025) reduces hypergraph generation to standard graph generation via a bipartite representation. However, no prior approach has explicitly integrated higher-order topology due to the stricter challenges of modeling multi-way rather than pairwise dependencies.
Appendix C Details of HOG-Diff
This section elucidates our spectral diffusion methodology, the parameterization of the score network, and the associated training and sampling procedures.
C.1 Overview
@@
As shown in Fig. 6, HOG-Diff employs a hierarchical, coarse-to-fine generation curriculum, where both forward diffusion and reverse denoising processes are decomposed into $K$ easy-to-learn subprocesses. Each subprocess is realized using the GOU bridge process.
@@
C.2 Spectral Diffusion
Classical graph diffusion approaches typically inject isotropic Gaussian noise directly into adjacency matrices $\bm{A}$, leading to various fundamental challenges. Firstly, the inherent non-uniqueness of graph representations implies that a graph with $n$ vertices can be equivalently modelled by up to $n!$ distinct adjacency matrices. This ambiguity requires a generative model that assigns probabilities uniformly across all equivalent adjacencies to accurately capture the graph’s inherent symmetry. Additionally, unlike densely distributed image data, graphs typically follow a Pareto distribution and exhibit sparsity (Ghavasieh & De Domenico, 2024), so that adjacency score functions lie on a low-dimensional manifold. Consequently, noise injected into out-of-support regions of the full adjacency space severely degrades the signal-to-noise ratio, impairing the training of the score-matching process. Even for densely connected graphs, isotropic noise distorts global message-passing patterns by encouraging message-passing on sparsely connected regions. Moreover, the adjacency matrix scales quadratically with the number of nodes, making the direct generation of adjacency matrices computationally prohibitive for large-scale graphs.
To address these challenges, inspired by Martinkus et al. (2022) and Luo et al. (2023a), we introduce noise in the eigenvalue domain of the graph Laplacian matrix $\bm{L}=\bm{D}-\bm{A}$, instead of the adjacency matrix $\bm{A}$, where $\bm{D}$ denotes the diagonal degree matrix. As a symmetric positive semi-definite matrix, the graph Laplacian can be diagonalized as $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{\top}$. Here, the orthogonal matrix $\bm{U}=[\bm{u}_{1},\cdots,\bm{u}_{n}]$ comprises the eigenvectors, and the diagonal matrix $\bm{\Lambda}=\operatorname{diag}(\lambda_{1},\cdots,\lambda_{n})$ holds the corresponding eigenvalues. The relationship between the Laplacian spectrum and the graph’s topology has been extensively explored (Chung, 1997). For instance, the low-frequency components of the spectrum capture the global structural properties such as connectivity and clustering, whereas the high-frequency components are crucial for reconstructing local connectivity patterns. Therefore, the target graph distribution $p(\bm{G}_{0})$ represents a joint distribution of $\bm{X}_{0}$ and $\bm{\Lambda}_{0}$, exploiting the permutation invariance and structural robustness of the Laplacian spectrum.
Consequently, we split the reverse-time SDE into two parts that share drift and diffusion coefficients as
$\left\{\begin{aligned} \mathrm{d}\bm{X}_{t}=&\left[\mathbf{f}_{k,t}(\bm{X}_{t})-g_{k,t}^{2}\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\right]\mathrm{d}\bar{t}+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}^{1}\\ \mathrm{d}\bm{\Lambda}_{t}=&\left[\mathbf{f}_{k,t}(\bm{\Lambda}_{t})-g_{k,t}^{2}\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\right]\mathrm{d}\bar{t}+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}^{2}\end{aligned}\right.$. (48)
Here, the superscript of $\bm{X}^{(k)}_{t}$ and $\bm{\Lambda}^{(k)}_{t}$ are dropped for simplicity, and $\mathbf{f}_{k,t}$ is determined according to Eq. 6.
In addition, we conduct a comparative evaluation of HOG-Diff under two generative settings: one operating directly in the adjacency matrix domain, and the other in the Laplacian spectral domain. Using a consistent hyperparameter search space, the results summarized in Tab. 6 show that generation in the spectral domain generally achieves comparable performance across most evaluation metrics. We adopt the Laplacian spectral domain as the default diffusion space in HOG-Diff, as the spectral approach is more efficient and better aligned with theoretical principles such as permutation invariance and signal concentration on low-dimensional manifolds.
C.3 Score Network Parametrization
The score network in HOG-Diff is a critical component responsible for estimating the score functions required to reverse the diffusion process effectively. The architecture of the proposed score network is depicted in Fig. 7. The input $\bm{A}_{t}$ is computed from $\bm{U}_{0}$ and $\bm{\Lambda}_{t}^{(k)}$ using the relation $\bm{A}_{t}=\bm{D}_{t}^{(k)}-\bm{L}^{(k)}_{t}$, where the Laplacian matrix is given by $\bm{L}^{(k)}_{t}=\bm{U}_{0}\bm{\Lambda}_{t}^{(k)}\bm{U}_{0}^{\top}$ and the diagonal degree matrix is given by $\bm{D}_{t}^{(k)}=\operatorname{diag}\left(\bm{L}_{t}^{(k)}\right)$. To enhance the input to the Attention module, we derive enriched node and edge features using the $l$-step random walk matrix obtained from the binarized $\bm{A}_{t}$. Specifically, the arrival probability vector is incorporated as additional node features, while the truncated shortest path distance derived from the same matrix is employed as edge features. Temporal information is integrated into the outputs of the Attention and GCN modules using Feature-wise Linear Modulation (FiLM) (Perez et al., 2018) layers, following sinusoidal position embeddings (Waswani et al., 2017).
Figure 7: Score Network Architecture of HOG-Diff. The score network integrates GCN and Attention blocks to capture both local and global features, and further incorporates time information through FiLM layers. These enriched outputs are subsequently concatenated and processed by separate feed-forward networks to produce predictions for $\nabla_{\bm{X}_{t}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}_{t}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, respectively.
A graph processing module is considered permutation invariant if its output remains unchanged under any permutation of its input, formally expressed as $f(\bm{G})=x\iff f(\pi(\bm{G}))=x$, where $\pi(\bm{G})$ represents a permutation of the input graph $\bm{G}$. It is permutation equivariant when the output undergoes the same permutation as the input, formally defined as $f(\pi(\bm{G}))=\pi(f(\bm{G}))$. It is worth noting that our score network model is permutation equivalent, as each model component avoids any node ordering-dependent operations.
C.4 Training and Sampling Procedure
The diffusion process in HOG-Diff is divided into $K$ hierarchical intervals, denoted by $\{[\tau_{k-1},\tau_{k}]\}_{k=1}^{K}$, where $0=\tau_{0}<\cdots<\tau_{k-1}<\tau_{k}<\cdots<\tau_{K}=T$. Within each interval $\tau_{k-1},\tau_{k}]$, we employ the GOU bridge process to ensure smooth transitions between intermediate states $\bm{G}_{\tau_{k-1}}$ and $\bm{G}_{\tau_{k}}$. We apply the cell complex filtering (CCF) operation at each interval to obtain structured, topologically meaningful intermediate states $\bm{G}_{\tau_{k}}:=\operatorname{CCF}(\bm{G},\mathcal{S},p)$. Specifically, CCF prunes nodes and edges that are not contained in the closure of any $p$-cell within a given cell complex $\mathcal{S}$. At the initial state, the filtering operation is defined as $\operatorname{CCF}(\bm{G},\mathcal{S},0)=\bm{G}$, i.e., the filtering operation leaves the input unchanged. A special case arises at the final step, where the intermediate state is initialized from Gaussian noise, i.e., $\operatorname{CCF}(\bm{G},\mathcal{S},K)\sim\mathcal{N}(\bm{0},\bm{I})$. Since the GOU bridge process naturally reduces to a standard diffusion process when the terminal distribution is Gaussian noise, we omit the GOU bridge in the final segment $\tau_{K-1},\tau_{K}]$, and instead use the Variance Preserving (VP) SDE (Ho et al., 2020; Song et al., 2021). In our experiments, we adopt a two-stage generation process, i.e., $K=2$. The intermediate state $\bm{G}_{\tau_{1}}$ is obtained via 2-cell complex filtering for molecule generation tasks, or via the 3-simplicial complex filtering for generic graph generation tasks. The rationale for this choice of filtering strategy is discussed in Sec. F.3.
To approximate the score functions $\nabla_{\bm{X}_{t}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}_{t}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, we employ a neural network $\bm{s}^{(k)}_{\bm{\theta}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$, as introduced in Sec. C.3. This model consists of a node ($\bm{s}^{(k)}_{\bm{\theta},\bm{X}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$) and a spectrum ($\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$) output. The network is trained by minimizing the following score-matching loss:
$\begin{split}\ell^{(k)}(\bm{\theta})=\mathbb{E}_{t,\bm{G}_{t},\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}}}\{\omega(t)[&c_{1}\|\bm{s}^{(k)}_{\bm{\theta},\bm{X}}-\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\|_{2}^{2}\\ +&c_{2}||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||_{2}^{2}]\},\end{split}$ (49)
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ controls the relative importance of vertices and spectrum.
The overall generation procedure is as follows. We sample $(\hat{\bm{X}}_{\tau_{K}},\hat{\bm{\Lambda}}_{\tau_{K}})$ from the prior distribution and select $\hat{\bm{U}}_{0}$ as an eigenbasis drawn from the training set. Reverse diffusion is then applied across multiple stages, sequentially generating $(\hat{\bm{X}}_{\tau_{K-1}},\hat{\bm{\Lambda}}_{\tau_{K-1}}),\dots,(\hat{\bm{X}}_{1},\hat{\bm{\Lambda}}_{1}),(\hat{\bm{X}}_{0},\hat{\bm{\Lambda}}_{0})$, where each stage is implemented via the diffusion bridge and initialized from the output of the previous step. Finally, plausible samples with higher-order structures can be reconstructed as $\hat{\bm{G}}_{0}=(\hat{\bm{X}}_{0},\hat{\bm{L}}_{0}=\hat{\bm{U}}_{0}\hat{\bm{\Lambda}}_{0}\hat{\bm{U}}_{0}^{\top})$.
We provide the pseudo-code of the training and sampling process in Alg. 1 and Alg. 2, respectively.